OpenStudy (anonymous):

Simplify -20 - (-5) (-2 squared) I got 0 , my dad -40 who is right and why ????

OpenStudy (anonymous):

this goes in the mathematics section if you need faster help btw

Directrix (directrix):

As written, -20 - (-5) (-2 squared) = -20 - (-5)*(-2²) = -20 - (-5)*(-4) = -20 - [(-5)*(-4)] = -20 - [20] = -20 - 20 = -40. The source of confusion is the difference between (-2)² and (-2²). The former equals 4 and the latter equals -4. Likewise, -2² = -4.